OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. B(0, 0), C(4, -4)

OpenStudy (cwrw238):

first find the slope of the line by dividing the difference in y-values by difference in x-values in this case that is (-4-0) / (4 - 0)

OpenStudy (cwrw238):

then use the general form y-y1 = m(x - x1) where m = slope you just found and (x1, y1) is one of the given points

OpenStudy (cwrw238):

in your case its easier to use (x1,y1) = (0,0)